OpenStudy (anonymous):

Find a ratio equivalent to each ratio. Then use the ratios to write a proportion ? 7/9

OpenStudy (jdoe0001):

well, if you multiply both numerator and denominator by a number, same top and bottom you'd get an equivalent fraction, and thus an equivalent ratio to write a proportion, I gather they just mean the NUMBER:NUMBER notation

Directrix (directrix):

@lbunton1 A proportion is two equal ratios written like this: a/b = c/d So, when you find a fraction that is equal to 7/9, then write the proportion as: 7/9 = ?/? where ?/? is the fraction that is equal to 7/9.